p^2q^2+x^2y^2= x^2q^2(x^2+y^2)


1
Expert's answer
2020-12-07T15:57:40-0500
"Solution"

To find the complete integral of the given PDE


The given PDE is


"p^2q^2+x^2y^2= x^2q^2(x^2+y^2)"

we may re-write it as


"p^2q^2+x^2y^2=x^2q^2(x^2+y^2)\\
or\\
p^2q^2+x^2y^2=x^4q^2+x^2y^2q^2\\
or\\
p^2q^2=x^4q^2+x^2y^2(q^2-1)\\
or\\
p^2=x^4+x^2y^2(1-\frac{1}{q^2})\\
or\\
p^2-x^4=x^2y^2(1-\frac{1}{q^2})\\
or\\
\frac{p^2}{x^2}-x^2=y^2(1-\frac{1}{q^2})\\
or\\
 \frac{p^2}{x^2}-x^2=y^2-\frac{y^2}{q^2}=a^2,\ say."

This is the special form "(f(p,x) =g(q,y)type)" , of the Charpit’s equations,


"\frac{p^2}{x^2}-x^2=a^2\ and\ y^2-\frac{y^2}{q^2}=a^2\\i.e\\
p=x\sqrt{x^2+a^2}\ and\ q=\frac{y}{\sqrt{y^2-a^2}}"

Putting these values of "p" and "q" in the equation


"dz=p\ dx+q\ dy"

We have


"dz=x\sqrt{x^2+a^2}\ dx+\frac{y}{\sqrt{y^2-a^2}}dy"

Integrating it, we get the complete integral as


"z=\frac13(x^2+a^2)^\frac32+\sqrt{y^2-a^2}+b"

Is the complete integral f the PDE